OpenStudy (anonymous):

yoo! i'm stuck help guys! solve and explain all steps in the following questions : \[\int\limits_{}\sqrt{sinx}dx\]

OpenStudy (anonymous):

-2 E(1/4 (pi-2 x)|2)+constant

OpenStudy (anonymous):

yeah i bet you are stuck. i don't think this has a nice closed form

OpenStudy (anonymous):

i took t=sqrt(sinx) so t^2 = sinx then i did it as cosxdx for dt but at the end step i stucekd ,that it being unintegrable

OpenStudy (anonymous):

forget it

OpenStudy (amistre64):

e^(-st) it :)

OpenStudy (amistre64):

euler would

OpenStudy (anonymous):

:( wolframalpha answers, but really, explain me the steps of it?

OpenStudy (amistre64):

if t^2 = sin(x); then 2t dt= cos(x) dx, and x = sin-1(t^2) \[\int \frac{2t^2}{cos(sin^{-1}(t^2))}dt\] not that that will be much easier, but if your gonna go that change of variable route it will turn to this i believe

OpenStudy (anonymous):

wolfram is telling you nothing. it is telling you that the integral is an integral

OpenStudy (anonymous):

well, i did the same thing, but after that part, i'm stuck...

OpenStudy (amistre64):



OpenStudy (amistre64):

\[\int \frac{2t^2}{\sqrt{(1-t^2)}}\ dt\]

OpenStudy (anonymous):

hmm.... so now?

OpenStudy (amistre64):

the 2t^2 should integrate by parts to 0 fairly quickly

OpenStudy (amistre64):

or at least get you to a step that is easier :) or not

OpenStudy (anonymous):

hmmm the triangle thing is really interesting and creative o.O

OpenStudy (amistre64):

yeah, we need the cos of an angle whose sin was t^2 :)

OpenStudy (anonymous):

its not able to do integrate by parts by the way... really did not deal a one like that before, its 2t^2 * dt / sqrt(1-t^2)

OpenStudy (amistre64):

lets rewrite and see if it gets better \[\int \frac{2t^2}{\sqrt{(1-t^2)}}\ dt\] \[\int 2t^2*\frac{1}{\sqrt{(1-t^2)}}\ dt\] \[=t *sin{-1}(t)-\int sin^{-1} (t)dt\] maybe?

OpenStudy (amistre64):

typoed it ... as usual

OpenStudy (anonymous):

hmm looks like good to me, did not see a mistake there

OpenStudy (amistre64):

\[=2t^2*sin^{-1}(t)-\int sin^{-1}(t)*4t\ dt\]

OpenStudy (anonymous):

that sin^-1*

OpenStudy (anonymous):

but here is the thing, how you defined cos(alfa) as equal to cos(sin^-1(t^2)) a trick on inverse functions?

OpenStudy (amistre64):

yes, its using the meaning of it all; cos(sin-1(t^2)) MEANS cos(angle), such that the sin(angle) = t^2

OpenStudy (anonymous):

ahh k, have to work on trig funcs a bit more...

OpenStudy (amistre64):

cos^2(a) + t^2 = 1 cos^2(a) = 1 - t^2 cos(a) = sqrt(1-t^2)

OpenStudy (anonymous):

k now i got it that part thx

OpenStudy (amistre64):

that parts a little deceptive since i did the math off

OpenStudy (amistre64):

sin(a) = t^2 ; sin^2(a) = t^4

OpenStudy (amistre64):

thats what i overlooked :)

OpenStudy (anonymous):

its sin^-1(t^2) yea not sin-1(t^2)

OpenStudy (anonymous):

:) :)

OpenStudy (anonymous):

if you google this the answers are pretty funny. most are just plain wrong, some people say "none" but a good response is here "So sorry, but this integral is not elementary. In fact, we can reduce it to an elliptic integral. To see this, let u = sin x, x = arcsin u, dx = du/√(1-u²) Then the integral becomes ∫√(u) du/√(1-u²), and, rationalising the denominator, ∫ √(u-u³ du/(1-u²) Since you have the square root of a cubic polynomial in the integrand, you now have an elliptic integral."

OpenStudy (anonymous):

it is not the case that just because you write down an integral you will be able to find some function in closed form whose derivative is the integrand.

OpenStudy (anonymous):

hmm...

OpenStudy (amistre64):

korcan never really does normal integrals :)

OpenStudy (anonymous):

in fact the truth is that if you pulled a function out of a (rather large imaginary) hat the probability that it would be integrable is 0